Calculate the number of liters of Hydrogen gas at STP that would be produced by the reaction of 6.20g of Magnesium with excess at HCl.

Mg+2HCl➡️MgCl2+H2
1
Expert's answer
2020-09-22T03:34:51-0400

As HCl is in excess, Mg will be the limiting reagent for this reaction.

n = m/M

M(Mg)=24.3 g/mol

n (Mg) = 6.20/24.3 = 0.3 mol

n (H2)=n (Mg) = 0.3 mol

n (H2) = V/22.4

V(H2) = n (H2) x 22.4 = 0.3 x 22.4 = 6.7 l